# Improving Group Robustness on Spurious Correlation Requires Preciser Group Inference

### v1

@@ 5 Discussion
In this work, we introduce GIC, a novel method for better inferring group labels and mitigating spurious correlations. Our analysis of misclassification cases reveals the semantic consistency phenomenon in GIC, which effectively disrupts spurious correlations and facilitates invariant learning.
Comparison data is essential for GIC and can be the labeled validation set, the subset of the unlabeled test set, or constructed non-uniformly from the training set, making GIC universally applicable. Appendix C.4 further emphasizes considering larger and cheaper unlabeled data as comparison data is necessary since an increased sample size of comparison data improves GIC’s performance. We also emphasize that the group distribution difference between comparison and training data can be subtle. Section 4.3 showcases GIC’s effectiveness on CelebA and Civilcomments, despite the high similarity in group distributions between the training and comparison data. In Appendix C.5, we further investigates the positive impact of group differences on GIC’s performance and propose readjusting the comparison data’s group distribution based on GIC’s inferred groups to further increase the distribution difference and the worst-group performance, as supported by experiments in Appendix C.5.
@@ Appendix A Proof Details
A.1 Proof of the transformation of $I(y^{tr};y^{tr}_{s,\mathbf{w}})$
Given $y^{tr}$, the maximization of mutual information $I(y^{tr};y^{tr}_{s,\mathbf{w}})$ can be transformed into the minimization of cross-entropy $H(y^{tr},y^{tr}_{s,\mathbf{w}})$.
@@ A.1 Proof of the transformation of $I(y^{tr};y^{tr}_{s,\mathbf{w}})$
Given $y^{tr}$, the correlation term is lower bounded by the difference between the entropy of $y^{tr}$ and the cross-entropy between $y^{tr}$ and $y^{tr}_{s,\mathbf{w}}$:
$I(y^{tr};y^{tr}_{s,\mathbf{w}})\geq H(y^{tr})-H(y^{tr},y^{tr}_{s,\mathbf{w}}).$
@@
First, we expand the mutual information term:
$\begin{split}I(y^{tr};y^{tr}_{s,\mathbf{w}})&=H(y^{tr})-H(y^{tr}|y^{tr}_{s,\mathbf{w}})\\ \end{split}$
@@
$\begin{split}H(y^{tr},y^{tr}_{s,\mathbf{w}})&=-\sum\mathbb{P}(y^{tr})\log\mathbb{P}(y^{tr}_{s,\mathbf{w}})\\ &=\sum\mathbb{P}(y^{tr})\log\frac{\mathbb{P}(y^{tr})}{\mathbb{P}(y^{tr}_{s,\mathbf{w}})}-\sum\mathbb{P}(y^{tr})\log\mathbb{P}(y^{tr})\\ &=\mathrm{KL}(y^{tr}||y^{tr}_{s,\mathbf{w}})+H(y^{tr})\\ &\geq\mathrm{KL}(y^{tr}||y^{tr}_{s,\mathbf{w}})+H(y^{tr}|y^{tr}_{s,\mathbf{w}})\\ &\geq H(y^{tr}|y^{tr}_{s,\mathbf{w}})\end{split}$
@@
$\begin{split}I(y^{tr};y^{tr}_{s,\mathbf{w}})&=H(y^{tr})-H(y|y^{tr}_{s,\mathbf{w}})\geq H(y^{tr})-H(y^{tr},y^{tr}_{s,\mathbf{w}})\end{split}$
According to Lemma A.1, when $y^{tr}$ is given, maximizing the mutual information $I(y^{tr};y^{tr}_{s,\mathbf{w}})$ can be equivalently transformed into minimizing $H(y^{tr},y^{tr}_{s,\mathbf{w}})$, which is more computationally tractable in practice.
A.2 Proof of Theorem 3.1
To prove Theorem 3.1, we first demonstrate the causal structure between $y_{s}$, $\mathbf{z}$, and $y$. Following the work [28], we assume $y_{s}\leftarrow\mathbf{z}\rightarrow y$. The fork causal structure [25] between $y_{s}$ , $\mathbf{z}$ and $y$ exhibits the following properties:
@@
According to the properties of KL divergence [11], we have the following:
$\begin{split}&\mathrm{KL}(\mathbb{P}(y^{tr},\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c},\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}}))-\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|y^{tr},y^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|y^{c},y^{c}_{s,\mathbf{w}}))\\ &=\mathbb{E}_{(y^{tr},\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}[\log\frac{\mathbb{P}(y^{tr},\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(y^{c},\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}})}]-\mathbb{E}_{(y^{tr},\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}[\log\frac{\mathbb{P}(\mathbf{z}^{tr}|y^{tr},y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(\mathbf{z}^{c}|y^{c},y^{c}_{s,\mathbf{w}})}]\\ &=\mathbb{E}_{(y^{tr},\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}[\log\frac{\mathbb{P}(y^{tr}|y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(y^{c}|y^{c}_{s,\mathbf{w}})}]\\ &=\mathbb{E}_{(y^{tr},y^{tr}_{s,\mathbf{w}})}[\log\frac{\mathbb{P}(y^{tr}|y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(y^{c}|y^{c}_{s,\mathbf{w}})}]\\ &=\mathrm{KL}(\mathbb{P}(y^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}|y^{c}_{s,\mathbf{w}}))\end{split}$ (14)
Term $\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|y^{tr},y^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|y^{c},y^{c}_{s,\mathbf{w}}))$ captures the similarity between the same group in the training and validation sets. Since $\mathbb{P}(\mathbf{z}^{tr}|y^{tr},y^{tr}_{s,\mathbf{w}})=\mathbb{P}(\mathbf{z}^{c}|y^{c},y^{c}_{s,\mathbf{w}})$, we have
$\begin{split}\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|y^{tr},y^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|y^{c},y^{c}_{s,\mathbf{w}}))&=\sum\mathbb{P}(y^{tr},y^{tr}_{s,\mathbf{w}})\sum\mathbb{P}(\mathbf{z}^{tr}|y^{tr},y^{tr}_{s,\mathbf{w}})\log\frac{\mathbb{P}(\mathbf{z}^{tr}|y^{tr},y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(\mathbf{z}^{c}|y^{c},y^{c}_{s,\mathbf{w}})}\\ &=\sum_{k}r^{tr}_{k}\sum P_{k}\log\frac{\mathbb{P}_{k}}{\mathbb{P}_{k}}=0\end{split}$ (15)
Consequently, maximizing the spurious term $\mathrm{KL}(\mathbb{P}(y^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}|y^{c}_{s,\mathbf{w}}))$ is tantamount to maximizing $\mathrm{KL}(\mathbb{P}(y^{tr},\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c},\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}}))$. ∎
@@
The tighter the bound is as $\mathrm{KL}(\mathbb{P}(y^{tr}|\mathbf{z}^{tr})||\mathbb{P}(y^{c}|\mathbf{z}^{c}))$ becomes smaller.
@@
Expanding the objective term, we have:
$\begin{split}&\mathrm{KL}(\mathbb{P}(y^{tr},\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c},\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}}))\\ &=\mathbb{E}_{(y^{tr},\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}[\log\frac{\mathbb{P}(y^{tr},\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(y^{c},\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}})}]\\ &=\mathbb{E}_{(y^{tr},\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}[\log\frac{\mathbb{P}(y^{tr},\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(y^{c},\mathbf{z}^{c},y^{c}_{s,\mathbf{w}})}]-\mathbb{E}_{y^{tr}_{s,\mathbf{w}}}[\log\frac{\mathbb{P}(y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(y^{c}_{s,\mathbf{w}})}]\\ &\overset{(a)}{=}\mathbb{E}_{(y^{tr},\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}[\log\frac{\mathbb{P}(y^{tr}|\mathbf{z}^{tr})\mathbb{P}(\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(y^{c}|\mathbf{z}^{c})\mathbb{P}(\mathbf{z}^{c},y^{c}_{s,\mathbf{w}})}]-\mathrm{KL}(\mathbb{P}(y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}_{s,\mathbf{w}}))\\ &=\mathbb{E}_{(y^{tr},\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}[\log\frac{\mathbb{P}(y^{tr}|\mathbf{z}^{tr})}{\mathbb{P}(y^{c}|\mathbf{z}^{c})}]+\mathbb{E}_{(y^{tr},\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}[\log\frac{\mathbb{P}(\mathbf{z}^{tr},y^{tr}_{s,\mathbf{w}})}{\mathbb{P}(\mathbf{z}^{c},y^{c}_{s,\mathbf{w}})}]-\mathrm{KL}(\mathbb{P}(y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}_{s,\mathbf{w}}))\\ &=\mathrm{KL}(\mathbb{P}(y^{tr}|\mathbf{z}^{tr})||\mathbb{P}(y^{c}|\mathbf{z}^{c}))+\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}}))\\ &\geq\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}}))\end{split}$ (17)
(a) achieves by the Property A.2 where we have:
$\begin{split}\mathbb{P}(y,\mathbf{z},y_{s})=\mathbb{P}(y|\mathbf{z})\mathbb{P}(y_{s}|\mathbf{z})\mathbb{P}(\mathbf{z})=\mathbb{P}(y|\mathbf{z})\mathbb{P}(y_{s},\mathbf{z})\end{split}$ (18)
By Equation 17, we prove the lower bound of $\mathrm{KL}(\mathbb{P}(y^{tr},\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c},\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}}))$ is $\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}}))$ and the tighter the bound is as $\mathrm{KL}(\mathbb{P}(y^{tr}|\mathbf{z}^{tr})||\mathbb{P}(y^{c}|\mathbf{z}^{c}))$ becomes smaller.
@@
$\mathrm{KL}(\mathbb{P}(y^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}|y^{c}_{s,\mathbf{w}}))\geq\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}}))$ (19)
@@
Based on Lemma A.3 and A.4, we can directly deduce:
$\mathrm{KL}(\mathbb{P}(y^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}|y^{c}_{s,\mathbf{w}}))\geq\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|y^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|y^{c}_{s,\mathbf{w}}))$ (20)

### v2

@@ 5 Discussion
In this work, we introduce GIC, a novel method for better inferring group labels and mitigating spurious correlations without any additional information requirement. Experiments on synthetic and real data reveal the enhancement of GIC in group label inference and its flexibility in integration with various invariant learning algorithms. Interestingly, our analysis of misclassification cases reveals the semantic consistency phenomenon in GIC, which effectively disrupts spurious correlations and facilitates invariant learning.
Comparison data is essential for GIC and can be the labeled validation set, the subset of the unlabeled test set, or constructed non-uniformly from the training set, making GIC universally applicable. Additional experiments in Appendix C.4 demonstrate the feasibility of constructing comparison data non-uniformly from the training set, achieving comparable performance on worst-group accuracy as directly using an validation set as comparison data. Moreover, Appendix C.4 further emphasizes that while non-uniform sampling from training data can construct comparison data, considering larger and cheaper unlabeled data as comparison data is necessary since an increased sample size of comparison data improves GIC’s performance. We also emphasize that the group distribution difference between comparison and training data can be subtle. Section 4.3 shows GIC’s effectiveness on CelebA and Civilcomments, despite the high similarity in group distributions between the training and comparison data. In Appendix C.6, we further investigate the positive impact of group differences on GIC’s performance. We propose readjusting the comparison data’s group distribution based on GIC’s inferred groups to further enhance the difference in group distribution and improve worst-group performance. This approach is supported by experiments detailed in Appendix C.6.
Acknowledgements
We would like to thank the anonymous reviewers and area chairs for their helpful comments. YH and DZ are supported by NSFC 62306252 and the central fund from HKU IDS.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Proof Details
Given $y^{tr}$, the maximization of mutual information $I(y^{tr};\hat{y}^{tr}_{s,\mathbf{w}})$ can be transformed into the minimization of cross-entropy $H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$.
@@
Given $y^{tr}$, the correlation term is lower bounded by the difference between the entropy of $y^{tr}$ and the cross-entropy between $y^{tr}$ and $\hat{y}^{tr}_{s,\mathbf{w}}$:
$I(y^{tr};\hat{y}^{tr}_{s,\mathbf{w}})\geq H(y^{tr})-H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}}).$
@@
For clearer expression, we omit the subscript $\mathbf{w}$ representing the training parameters in $\hat{y}_{s,\mathbf{w}}$. We first expand the mutual information term:
$\begin{split}I(y^{tr};\hat{y}^{tr}_{s})&=H(y^{tr})-H(y^{tr}|\hat{y}^{tr}_{s})\\ \end{split}$
@@
$\begin{split}H(y^{tr},\hat{y}^{tr}_{s})&=-\sum\mathbb{P}(y^{tr})\log\mathbb{P}(\hat{y}^{tr}_{s})\\ &=\sum\mathbb{P}(y^{tr})\log\frac{\mathbb{P}(y^{tr})}{\mathbb{P}(\hat{y}^{tr}_{s})}-\sum\mathbb{P}(y^{tr})\log\mathbb{P}(y^{tr})\\ &=\mathrm{KL}(y^{tr}||\hat{y}^{tr}_{s})+H(y^{tr})\\ &\geq\mathrm{KL}(y^{tr}||\hat{y}^{tr}_{s})+H(y^{tr}|\hat{y}^{tr}_{s})\\ &\geq H(y^{tr}|\hat{y}^{tr}_{s}).\end{split}$
@@
$\begin{split}I(y^{tr};\hat{y}^{tr}_{s})&=H(y^{tr})-H(y|\hat{y}^{tr}_{s})\geq H(y^{tr})-H(y^{tr},\hat{y}^{tr}_{s}).\end{split}$
According to Lemma A.1, when $y^{tr}$ is given, maximizing the mutual information $I(y^{tr};\hat{y}^{tr}_{s,\mathbf{w}})$ can be transformed into minimizing $H(y^{tr},\hat{y}^{tr}_{s,\mathbf{w}})$, which is more computationally tractable in practice. ∎
A.1 Proof of Theorem 3.1
To prove Theorem 3.1, we first demonstrate the causal structure between $y_{s}$, $\mathbf{z}$, and $y$. Following the work (Li et al., 2022), we assume $y_{s}\leftarrow\mathbf{z}\rightarrow y$. The fork causal structure (Lagnado & Sloman, 2019) between $y_{s}$ , $\mathbf{z}$ and $y$ exhibits the following properties:
@@
Naturally, the proxy $\hat{y}_{s,\mathbf{w}}$ for the true spurious attribute label $y_{s}$ should also satisfy the aforementioned causal properties.
@@
As stated in Section 3.2, different variable names are used to emphasize various data sources. In the proof section, we consider to a more formal expression for clarity. That is, we consider different distribution notations with the same variable names and also ignore the subscript $\mathbf{w}$.
$\begin{split}&\mathbb{P}_{tr}(y|\hat{y}_{s}):=\mathbb{P}(y^{\mathrm{tr}}|\hat{y}^{tr}_{s})\\ &\mathbb{P}_{\mathrm{c}}(y|\hat{y}_{s}):=\mathbb{P}(y^{c}|\hat{y}^{c}_{s}).\end{split}$ (15)
According to the properties of KL divergence (Cover, 1999), we have the following:
$\begin{split}&\mathrm{KL}(\mathbb{P}_{tr}(y,\mathbf{z}|\hat{y}_{s})||\mathbb{P}_{c}(y,\mathbf{z}|\hat{y}_{s}))-\mathrm{KL}(\mathbb{P}_{tr}(\mathbf{z}|y,\hat{y}_{s})||\mathbb{P}_{c}(\mathbf{z}|y,\hat{y}_{s}))\\ &=\mathbb{E}_{(y,\mathbf{z},\hat{y}_{s})}\Big{[}\log\frac{\mathbb{P}_{tr}(y,\mathbf{z}|\hat{y}_{s})}{\mathbb{P}_{c}(y,\mathbf{z}|\hat{y}_{s})}\Big{]}-\mathbb{E}_{(y,\mathbf{z},\hat{y}_{s})}\Big{[}\log\frac{\mathbb{P}_{tr}(\mathbf{z}|y,\hat{y}_{s})}{\mathbb{P}_{c}(\mathbf{z}|y,\hat{y}_{s})}\Big{]}\\ &=\mathbb{E}_{(y,\mathbf{z},\hat{y}_{s})}\Big{[}\log\frac{\mathbb{P}_{tr}(y|\hat{y}_{s})}{\mathbb{P}_{c}(y|\hat{y}_{s})}\Big{]}\\ &=\mathbb{E}_{(y,\hat{y}_{s})}\Big{[}\log\frac{\mathbb{P}_{tr}(y|\hat{y}_{s})}{\mathbb{P}_{c}(y|\hat{y}_{s})}\Big{]}\\ &=\mathrm{KL}(\mathbb{P}_{tr}(y|\hat{y}_{s})||\mathbb{P}_{c}(y|\hat{y}_{s})).\end{split}$ (16)
Since term $\mathrm{KL}(\mathbb{P}_{tr}(\mathbf{z}|y,\hat{y}_{s})||\mathbb{P}_{c}(\mathbf{z}|y,\hat{y}_{s}))$ captures the similarity between the same group in the training and validation sets, i.e., $\mathbb{P}_{tr}(\mathbf{z}|y,\hat{y}_{s})=\mathbb{P}_{c}(\mathbf{z}|y,y_{s})$, we have
$\begin{split}\mathrm{KL}(\mathbb{P}_{tr}(\mathbf{z}|y,\hat{y}_{s})||\mathbb{P}_{c}(\mathbf{z}|y,\hat{y}_{s}))=\sum_{g\in\mathcal{G}}\mathbb{P}_{tr}((y,\hat{y}_{s})=g)\sum_{z}\mathbb{P}_{tr}(\mathbf{z}=z|(y,\hat{y}_{s})=g)\log\frac{\mathbb{P}_{tr}(\mathbf{z}=z|(y,\hat{y}_{s})=g)}{\mathbb{P}_{c}(\mathbf{z}=z|(y,y_{s})=g)}=0.\end{split}$ (17)
Therefore, maximizing the spurious term $\mathrm{KL}(\mathbb{P}_{tr}(y|\hat{y}_{s})||\mathbb{P}_{c}(y|\hat{y}_{s}))$ is tantamount to maximizing $\mathrm{KL}(\mathbb{P}_{tr}(y,\mathbf{z}|\hat{y}_{s})||\mathbb{P}_{c}(y,\mathbf{z}|\hat{y}_{s}))$. ∎
@@
We express our variables in the same manner as in Lemma A.3 and omit the subscript $\mathbf{w}$.
$\begin{split}&\mathrm{KL}(\mathbb{P}_{tr}(y,\mathbf{z}|\hat{y}_{s})||\mathbb{P}_{c}(y,\mathbf{z}|\hat{y}_{s}))\\ &=\mathbb{E}_{(y,\mathbf{z},\hat{y}_{s})}\Big{[}\log\frac{\mathbb{P}_{tr}(y,\mathbf{z}|\hat{y}_{s})}{\mathbb{P}_{c}(y,\mathbf{z}|\hat{y}_{s})}\Big{]}\\ &=\mathbb{E}_{(y,\mathbf{z},\hat{y}_{s})}\Big{[}\log\frac{\mathbb{P}{tr}(y,\mathbf{z},\hat{y}_{s})}{\mathbb{P}_{c}(y,\mathbf{z},y_{s})}\Big{]}-\mathbb{E}_{\hat{y}_{s}}\Big{[}\log\frac{\mathbb{P}_{tr}(\hat{y}_{s})}{\mathbb{P}_{c}(\hat{y}_{s})}\Big{]}.\end{split}$ (19)
By the Property A.2, we derive:
$\begin{split}\mathbb{P}(y,\mathbf{z},\hat{y}_{s})=\mathbb{P}(y|\mathbf{z})\mathbb{P}(\hat{y}_{s}|\mathbf{z})\mathbb{P}(\mathbf{z})=\mathbb{P}(y|\mathbf{z})\mathbb{P}(\hat{y}_{s},\mathbf{z}).\end{split}$ (20)
Hence, Equation (19) can be simplified as follows
$\begin{split}&\mathrm{KL}(\mathbb{P}_{tr}(y,\mathbf{z}|\hat{y}_{s})||\mathbb{P}_{c}(y,\mathbf{z}|\hat{y}_{s}))\\ &=\mathbb{E}_{(y,\mathbf{z},\hat{y}_{s})}\Big{[}\log\frac{\mathbb{P}_{tr}(y|\mathbf{z})\mathbb{P}_{tr}(\mathbf{z},\hat{y}_{s})}{\mathbb{P}_{c}(y|\mathbf{z})\mathbb{P}_{c}(\mathbf{z},\hat{y}_{s})}\Big{]}-\mathrm{KL}(\mathbb{P}_{tr}(\hat{y}_{s})||\mathbb{P}_{c}(\hat{y}_{s}))\\ &=\mathbb{E}_{(y,\mathbf{z},\hat{y}_{s})}\Big{[}\log\frac{\mathbb{P}_{tr}(y|\mathbf{z})}{\mathbb{P}_{c}(y|\mathbf{z})}\Big{]}+\mathbb{E}_{(y,\mathbf{z},\hat{y}_{s})}\Big{[}\log\frac{\mathbb{P}_{tr}(\mathbf{z},\hat{y}_{s})}{\mathbb{P}_{c}(\mathbf{z},\hat{y}_{s})}\Big{]}-\mathrm{KL}(\mathbb{P}_{tr}(y_{s})||\mathbb{P}_{c}(\hat{y}_{s}))\\ &=\mathrm{KL}(\mathbb{P}_{tr}(y|\mathbf{z})||\mathbb{P}_{c}(y|\mathbf{z}))+\mathrm{KL}(\mathbb{P}_{tr}(\mathbf{z}|\hat{y}_{s})||\mathbb{P}_{c}(\mathbf{z}|\hat{y}_{s}))\\ &\geq\mathrm{KL}(\mathbb{P}_{tr}\Big{(}\mathbf{z}|\hat{y}_{s})||\mathbb{P}_{c}(\mathbf{z}|\hat{y}_{s})\Big{)}.\end{split}$ (21)
By Equation 21, we prove the lower bound of $\mathrm{KL}(\mathbb{P}_{tr}(y,\mathbf{z}|y_{s})||\mathbb{P}_{c}(y,\mathbf{z}|y_{s}))$ is $\mathrm{KL}(\mathbb{P}_{tr}(\mathbf{z}|y_{s})||\mathbb{P}_{c}(\mathbf{z}|y_{s,\mathbf{w}}))$. The proof is complete. ∎
@@
$\mathrm{KL}(\mathbb{P}(y^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}|\hat{y}^{c}_{s,\mathbf{w}}))\geq\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|\hat{y}^{c}_{s,\mathbf{w}}))$ (22)
@@
Based on Lemma A.3 and Lemma A.4, we can directly deduce:
$\mathrm{KL}(\mathbb{P}(y^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(y^{c}|\hat{y}^{c}_{s,\mathbf{w}}))\geq\mathrm{KL}(\mathbb{P}(\mathbf{z}^{tr}|\hat{y}^{tr}_{s,\mathbf{w}})||\mathbb{P}(\mathbf{z}^{c}|\hat{y}^{c}_{s,\mathbf{w}})).$ (23)